### v2

@@ Abstract
The idea to create Global Benchmark Database (GBD) arose from the need to make available and sustainably maintain benchmark instances and their metadata for the problem of propositional satisfiability (SAT). To this end, we specified a simple hash function for instance identification to serve as the primary key for SAT instance data sets. A proof of concept was presented and discussed at the 2018 Pragmatics of SAT (POS) workshop, demonstrating the specification of a SAT instance identifier and methods for labeling and querying for instances [15]. From there, GBD has matured into a comprehensive suite of tools for provisioning and sustainably maintaining benchmark instances of various problem domains and their metadata. Supported problem domains include propositional satisfiability (SAT), maximum satisfiability (MaxSAT), and pseudo-Boolean optimization (PBO).
In essence, the purpose of GBD is to act as a conduit between data science and empirical research on NP-hard problem classes by facilitating the seamless integration of benchmark data into existing workflows. GBD provides tools for distributing benchmark instances and feature databases, as well as tools for transforming instances and extracting instance features. It also provides a set of prebuilt feature databases for identifying instance equivalence classes, categories, or labels, as well as analytical instance features. Examples of successful applications of GBD include the sanitization and selection of benchmarks for SAT competitions [13], domain-specific solver evaluations [8], and the analysis of solver portfolios and solver prediction models [2]. Moreover, the authors of the latest award-winning state-of-the-art SAT solvers use GBD in their empirical evaluations. [11, 21]. GBD is open source and available under the MIT license [14].
@@ 2 Data Model and Queries
Given the context of cnf instances of the SAT problem and the context kis of graph-instances of the -Independent Set problem. Any SAT instance can be transformed to a graph and a number in such a way that the satisfiability of the SAT instance is equivalent to the existence of an independent set of size in the graph [19]. A context mapping feature can be used to map instances in cnf to their corresponding instances in kis.
@@ 3.3 Web Interface
The GBD web interface is implemented as a RESTful web service [7] that provides access to the benchmarks and databases of the environment it runs in, and can be started with the command gbd serve. While the GBD command-line tool operates independently of the web service, the web service uses the same GBD API and database configuration as the command-line tool. Our instance of the GBD web interface, accessible via {https://benchmark-database.de}, is configured to run behind an Nginx reverse proxy [20] and is hosted in a Docker container [17]. At the time of writing, we provide access to more than 100 000 benchmark instances from the cnf, wcnf, and opb contexts, with prebuilt feature databases for them.
@@ 3.4 Python Interface
The GBD Python API is wrapped in the class gbd_core.api.GBD and its documentation can be found at https://udopia.github.io/gbd. With the Python API, GBD data sources can be directly integrated into Python scripts and Jupyter notebooks to be used in evaluation and analysis. Use case LABEL:cmd:gbd:api illustrates the use of the Python API to query for instances and features. Line 2 creates a GBD object that connects to the GBD data sources base and meta, which are passed to the constructor as a list of paths. Line 3 creates a list of feature names from the base data source and adds the feature family from the meta data source. Line 4 queries for the given features of cnf instances from the Main track of SAT Competition 2023. The result is returned as a Pandas [16] DataFrame for further analysis.555We provide illustrative examples of benchmark data analyses, including portfolio analysis, category prediction, and category-specific ranking, accessible via https://udopia.github.io/gbdeval/.
@@ 5 Related and Future Work
SatLib was the first public collection of benchmark instances in the SAT problem domain [12], and similar collections have been created for other problem domains such as MaxSAT [1], Quantified Boolean Formulas (QBF) [9], SAT Modulo Theories (SMT) [4], and Mixed Integer Programming (MIP) [10]. SatEx is the first web-based framework for reproducible execution and evaluation of SAT solver experiments [22], followed by the EDACC framework [3]. Instance features were used to predict the fastest solver for an instance [23] and to reduce redundancy in experiments [18]. Aslib is a library of data sets for training and evaluating solver prediction models [5]. Meta-features such as the instance domain have been used to study specialized heuristic configurations [6].

### v3

@@ Abstract
The idea to create Global Benchmark Database (GBD) arose from the need to make available and sustainably maintain benchmark instances and their metadata for the problem of propositional satisfiability (SAT). To this end, we specified a simple hash function for instance identification to serve as the primary key for SAT instance data sets. A proof of concept was presented and discussed at the 2018 Pragmatics of SAT (POS) workshop, demonstrating the specification of a SAT instance identifier and methods for labeling and querying for instances [Iser:2018:GBD]. From there, GBD has matured into a comprehensive suite of tools for provisioning and sustainably maintaining benchmark instances of various problem domains and their metadata. Supported problem domains include propositional satisfiability (SAT), maximum satisfiability (MaxSAT), and pseudo-Boolean optimization (PBO).
In essence, the purpose of GBD is to act as a conduit between data science and empirical research on NP-hard problem classes by facilitating the seamless integration of benchmark data into existing workflows. GBD provides tools for distributing benchmark instances and feature databases, as well as tools for transforming instances and extracting instance features. It also provides a set of prebuilt feature databases for identifying instance equivalence classes, categories, or labels, as well as analytical instance features. Examples of successful applications of GBD include the sanitization and selection of benchmarks for SAT competitions [Iser:SC2023Bench], domain-specific solver evaluations [SC2020:AIJ], and the analysis of solver portfolios and solver prediction models [Bach:2022:kPortfolios]. Moreover, the authors of the latest award-winning state-of-the-art SAT solvers use GBD in their empirical evaluations. [Haberlandt:2023:SBVA, Schreiber:2023:Diss]. GBD is open source and available under the MIT license [Iser_GBD_Tools_2023].
@@ 2 Data Model and Queries
Given the context of cnf instances of the SAT problem and the context kis of graph-instances of the -Independent Set problem. Any SAT instance can be transformed to a graph and a number in such a way that the satisfiability of the SAT instance is equivalent to the existence of an independent set of size in the graph [Papadimitriou:1994:Complexity]. A context mapping feature can be used to map instances in cnf to their corresponding instances in kis.
@@ 3.3 Web Interface
The GBD web interface is implemented as a RESTful web service [Fielding:2000:REST] that provides access to the benchmarks and databases of the environment it runs in, and can be started with the command gbd serve. While the GBD command-line tool operates independently of the web service, the web service uses the same GBD API and database configuration as the command-line tool. Our instance of the GBD web interface, accessible via {https://benchmark-database.de}, is configured to run behind an Nginx reverse proxy [Reese:2008:Nginx] and is hosted in a Docker container [Merkel:2014:Docker]. At the time of writing, we provide access to more than 100 000 benchmark instances from the cnf, wcnf, and opb contexts, with prebuilt feature databases for them.
@@ 3.4 Python Interface
The GBD Python API is wrapped in the class gbd_core.api.GBD and its documentation can be found at https://udopia.github.io/gbd. With the Python API, GBD data sources can be directly integrated into Python scripts and Jupyter notebooks to be used in evaluation and analysis. Use case LABEL:cmd:gbd:api illustrates the use of the Python API to query for instances and features. Line 2 creates a GBD object that connects to the GBD data sources base and meta, which are passed to the constructor as a list of paths. Line 3 creates a list of feature names from the base data source and adds the feature family from the meta data source. Line 4 queries for the given features of cnf instances from the Main track of SAT Competition 2023. The result is returned as a Pandas [McKinney:2010:Pandas] DataFrame for further analysis.555We provide illustrative examples of benchmark data analyses, including portfolio analysis, category prediction, and category-specific ranking, accessible via https://udopia.github.io/gbdeval/.
@@ 5 Related and Future Work
SatLib was the first public collection of benchmark instances in the SAT problem domain [Hoos:2000:SatLib], and similar collections have been created for other problem domains such as MaxSAT [maxsatlib], Quantified Boolean Formulas (QBF) [qbflib], SAT Modulo Theories (SMT) [smtlib], and Mixed Integer Programming (MIP) [miplib]. SatEx is the first web-based framework for reproducible execution and evaluation of SAT solver experiments [Simon:2001:SatEx], followed by the EDACC framework [Balint:2010:EDACC]. Instance features were used to predict the fastest solver for an instance [Hoos:2008:SATzilla] and to reduce redundancy in experiments [Manthey:2016:Structure]. Aslib is a library of data sets for training and evaluating solver prediction models [Hoos:2016:Aslib]. Meta-features such as the instance domain have been used to study specialized heuristic configurations [Elffers:2018:Heuristics].